# Latent Denoising Makes Good Visual Tokenizers

### v1

@@ 5 Experiments
We now empirically analyze our proposed tokenizer, beginning with a series of ablation studies on noising strategies, followed by comprehensive comparisons across various generative modeling frameworks. Lastly, we benchmark against leading systems.
@@ 5.1 Main Properties
We study the generalizability of a tokenizer across non-autoregressive (non-AR) and autoregressive (AR) generative paradigms. To this end, we use SiT-B [42] and MAR-B [36] as representative non-AR and AR models, respectively. For all ablation studies, we adopt small encoders and base-sized decoders in our tokenizers.
@@ 5.1.1 Properties of Latent Noising
Figure 2: Ablation on latent noise design. (a) Additive vs. interpolative noise. Interpolative noise clearly outperforms additive noise for both MAR [36] and SiT [42]. Both interpolative and additive latent noise lead to improved performance for MAR. (b) Latent noise standard deviation ($\gamma$). Our $l$-DeTok remains robust across various noise standard deviations. Generally, increasing $\gamma$ improves generation quality, with best results achieved around $\gamma=3.0$.
Interpolative vs. additive noise.
An important design choice in our $l$-DeTok is the use of interpolative latent noise instead of additive noise, which we ablate here. Specifically, we compare two latent noising variants: interpolative noise, i.e., $\mathbf{x}^{\prime}=(1-\tau)\mathbf{x}+\tau\boldsymbol{\varepsilon}$ (Eq. 4), and additive noise, i.e., $\mathbf{x}^{\prime}=\mathbf{x}+\tau\boldsymbol{\varepsilon}$. We set the noise standard deviation to $\gamma=1.0$ here. Figure 2-(a) presents the results.
Interpolative noise clearly outperforms additive noise for both SiT [42] and MAR [36]. This aligns with our expectation: interpolative noise explicitly ensures latent embeddings are heavily corrupted at high noise levels. In contrast, additive noise can potentially create shortcuts by making original signals remain dominant, reducing the effectiveness of the additive noise. Nonetheless, we observe the additive latent noise still improves generative performance for MAR but not for SiT. Understanding the underlying factors responsible for these model-specific differences would be an intriguing future direction, e.g., joint vs. factorized latent distribution estimation.
Noise standard deviation.
Figure 2-(b) studies the effect of noise standard deviation ($\gamma$ in Eq. 4). Both SiT and MAR consistently improve with interpolative latent noise across all tested levels. Performance peaks at moderately high standard deviation, indicating that stronger corruption generally yields more effective latents. This result confirms our key hypothesis: challenging denoising tasks naturally produce robust, downstream-aligned latents that benefit generative modeling.
@@ 5.1.2 Properties of Masking
Masking ratio.
We examine how varying the maximal masking ratio $M$ (Eq. 5) influences generation quality. As shown in Figure 3-(a), both SiT and MAR benefit from masking-based tokenizer training in generation quality. Masking ratios between 70% and 90% consistently yield stronger performance compared to low masking ratios (e.g., 30%), favoring high degrees of masking. This behavior mirrors observations from latent denoising, i.e., challenging denoising is more beneficial, indicating a common underlying principle under the deconstruction-reconstruction strategy.
Interestingly, SiT [42] still benefits from our masking-based $l$-DeTok despite not explicitly handling masked inputs during its own training. We hypothesize that masking implicitly promotes encoders to learn embeddings that are inherently robust other noise types, such as diffusion-based perturbations.
Constant vs. randomized masking ratio.
Figure 3-(b) compares randomized masking ratios against constant ones used in MAE [30]. For constant masking ratio experiments, we fine-tune the tokenizer decoder for an additional 10 epochs on full-visible latents to mitigate the distribution mismatch between training and inference, since fully visible inputs are absent during constant-ratio training. From Figure 3-(b), we see that randomized masking outperforms constant masking. Randomized masking encourages latent embeddings that are robust to varying corruption levels. This is more aligned with the downstream tasks, i.e., denoising across diverse levels of corruption.
[…]
@@ 5.1.3 Joint Denoising
Prior ablations indicate that both latent noising and masking can independently improve generation quality, with latent noising showing a stronger effect. Here, we investigate the effect of joint denoising. Based on prior results, we fix the noise standard deviation to $\gamma=3.0$ and the masking ratio to $M=0.7$. Figure 5 and Table 5 summarize the results.
Figure 4: Impact of tokenizer training strategies on generative performance. We compare our baseline tokenizer with our $l$-DeTok variants: masking-only ($M{=}0.7$), latent noising-only ($\gamma{=}3.0$), and their combination (joint noising). Both masking and latent noising independently improve generation quality, with latent noising showing a stronger effect. Joint noising further improves performance for MAR, particularly in inception scores (IS), but provides limited additional benefit for SiT when latent noising is already applied. FID@50k scores are detailed in Tab. 5.
@@
With joint denoising, our $l$-DeTok achieves FID scores of 5.50 (SiT-B) and 2.65 (MAR-B) with CFG (Table 5). In comparison, our baseline tokenizer—trained with identical settings but without any noise—obtains significantly worse results: 6.97 (SiT-B) and 3.31 (MAR-B). We observe that joint denoising is more effective for MAR, further pushing both FID and IS, but provides limited additional benefit for SiT when latent noising is already applied. This indicates that latent denoising is essential, while the masking-based denoising can be optional.
Lastly, with joint denoising, we increase the encoder to base size, train for 200 epochs, and enable the GAN loss starting from epoch 100 (+Extended in Table 5). Under this setting, performance improves to 5.13 (SiT) and 2.43 (MAR). We adopt this improved tokenizer for all subsequent evaluations.
@@ 5.2 Generalization Experiments
To comprehensively evaluate tokenizer generalizability, we compare performance across six representative generative models—three non-autoregressive (DiT [46], SiT [42], LightningDiT [64]) and three autoregressive (MAR [36], RandomAR, RasterAR)—using publicly available tokenizers introduced before. RandomAR and RasterAR are Transformer-based decoder-only (causal attention) variants adapted from RAR [68]. We modify them to support decoding continuous tokens via diffloss MLPs [36]. RandomAR generates tokens in random order while RasterAR uses a raster-scan order. We use base-sized models and train them for 100 epochs for experiments here.
Comparisons with standard convolutional tokenizers.
[…]
Comparisons with semantics-distilled tokenizers.
Our $l$-DeTok generalizes significantly better than semantics-distilled tokenizers. Table 1 also compares our method against recent semantics-distilled tokenizers such as VA-VAE [64] and MAETok [6], which distill semantics features from powerful pretrained encoders [43, 49]. Surprisingly, we empirically find that these tokenizers, despite their promising performance for non-AR models, do not generalize well to AR models. Specifically, FID scores (with CFG) degrade dramatically from 3.71, 11.78, and 7.99 to 16.66, 38.13, and 15.88 for MAR, RandomAR, and RasterAR, respectively. Previous studies implicitly presume that tokenizer improvements observed in one generative paradigm naturally extend to others. Yet, our experiments challenge this assumption, revealing a previously unrecognized gap in tokenizer transferability: tokenizer effectiveness in one generative paradigm does not necessarily transfer to others.
In sharp contrast, our method generalizes significantly better across both non-AR and AR models. Importantly, our $l$-DeTok requires no semantics distillation from large-scale pretrained encoders yet considerably surpasses standard tokenizer competitors. This lifts external dependencies and highlights a new and previously overlooked direction in tokenizer research.
@@
VA-VAE [64] 0.28 16.66 144.5 38.13 68.3 15.88 160.5 4.33 221.1 4.91 213.9 2.86 275.1
MAETok [6] 0.48 6.99 201.8 24.83 97.6 15.92 127.2 4.77 243.2 5.24 224.7 3.92 273.3
@@ 5.3 Benchmarking with Previous Systems
We compare against leading generative systems in Table 2. For this experiment, we train MAR-B and MAR-L for 800 epochs. Simply adopting our tokenizer, without altering the MAR architecture, substantially improves the generative performance: MAR-B achieves an FID of 1.55 (from 2.31), and MAR-L further improves to 1.35 (from 1.78). Notably, our MAR-B and MAR-L both match or surpass the previously best-performing huge-size MAR model (1.35 vs. 1.55). Qualitative results are provided in Figure 6.
Table 2: System-level comparison on ImageNet 256$\times$256 class-conditioned generation. Our approach enables MAR models [36] to achieve leading results without relying on semantics distillation. : With additional decoder fine-tuning (see Sec. A.1 for details).
@@
SiT-XL + REPA [70] 675M 5.90 157.8 0.70 0.69 1.42 305.7 0.80 0.64
SiT-XL + MAETok [6] 675M 2.31 216.5 - - 1.67 311.2 - -
LightningDiT + MAETok [6] 675M 2.21 208.3 - - 1.73 308.4 - -
LightningDiT + VAVAE [64] 675M 2.17 205.6 0.77 0.65 1.35 295.3 0.79 0.65
DDT-XL [63] 675M 6.27 154.7 0.68 0.69 1.26 310.6 0.79 0.65
@@
ADM [16] 554M 10.94 101.0 0.69 0.63 4.59 186.7 0.82 0.52
VDM+++++ + [32] 2B 2.40 225.3 - - 2.12 267.7 - -
MaskGIT [5] 227M 6.18 182.1 0.80 0.51 - - - -
MAGVIT-v2 [67] 307M 3.65 200.5 - - 1.78 319.4 - -
LDM-4 [51] 400M 10.56 103.5 0.71 0.62 3.60 247.7 0.87 0.48
DiT-XL/2 [45] 675M 9.62 121.5 0.67 0.67 2.27 278.2 0.83 0.57
SiT-XL/2 [42] 675M 8.30 - - - 2.06 270.3 0.82 0.59
MDTv2-XL/2 [23] 676M 5.06 155.6 0.72 0.66 1.58 314.7 0.79 0.65
MaskDiT [23] 675M 5.69 178.0 0.74 0.60 2.28 276.6 0.89 0.61
VAR-d30 [59] 2.0B - - - - 1.92 323.1 0.82 0.59
LlamaGen-3B  [58] 3.1B - - - - 2.18 263.3 0.81 0.58
RandAR-XXL  [44] 1.4B - - - - 2.15 322.0 0.79 0.62
FlowAR-H [50] 1.9B - - - - 1.65 296.5 0.83 0.60
CausalFusion [14] 676M 3.61 180.9 0.75 0.66 1.77 282.3 0.82 0.61
MAR-B + MAR-VAE [36] 208M 3.48 192.4 0.78 0.58 2.31 281.7 0.82 0.57
MAR-L + MAR-VAE [36] 479M 2.60 221.4 0.79 0.60 1.78 296.0 0.81 0.60
MAR-H + MAR-VAE [36] 943M 2.35 227.8 0.79 0.62 1.55 303.7 0.81 0.62
MAR-B [36] + our $l$-DeTok 208M 2.79 195.9 0.80 0.60 1.61 289.7 0.81 0.62
MAR-B [36] + our l-DeTok 208M 2.94 195.5 0.80 0.59 1.55 291.0 0.81 0.62
MAR-L [36] + our $l$-DeTok 479M 1.84 238.4 0.82 0.60 1.43 303.5 0.82 0.61
MAR-L [36] + our l-DeTok 479M 1.86 238.6 0.82 0.61 1.35 304.1 0.81 0.62

### v2

@@ 5.1 Main Properties
We use SiT-B (Ma et al., 2024) and MAR-B (Li et al., 2024a) as representative non-autoregressive (non-AR) and autoregressive (AR) generative models to study the generalizability of a tokenizer. Quantitative reconstruction results for tokenizers trained under different noises are in Sec. B.1.
[…]
@@ 5.1.1 Properties of Latent Noising
Interpolative vs. additive noise. An important design choice in our $l$-DeTok is the use of interpolative latent noise instead of additive noise, which we ablate here. Specifically, we compare two latent noising variants: interpolative noise, i.e., $\mathbf{x}^{\prime}=(1-\tau)\mathbf{x}+\tau\bm{\varepsilon}$ (Eq. 4), and additive noise, i.e., $\mathbf{x}^{\prime}=\mathbf{x}+\tau\bm{\varepsilon}$. We set the noise standard deviation to $\gamma=1.0$ here. Figure 2-(a) presents the results: Interpolative noise clearly outperforms additive noise for both SiT (Ma et al., 2024) and MAR (Li et al., 2024a). This aligns with our expectation: interpolative noise ensures latent embeddings are heavily corrupted at high noise levels. In contrast, additive noise can potentially create shortcuts by making original signals remain dominant, reducing the effectiveness of the additive noise. Nonetheless, we observe the additive latent noise still improves MAR but not SiT.
Noise standard deviation. Figure 2-(b) studies the effect of noise standard deviation ($\gamma$ in Eq. 4). Both SiT and MAR consistently improve with interpolative latent noise across all tested levels. Performance peaks at moderately high standard deviation, indicating that stronger corruption generally yields more effective latents. This result confirms our key hypothesis: challenging denoising tasks naturally produce robust, downstream-aligned latents that benefit generative modeling.
[…]
MAR-B SiT-B
$\tau=0$ (baseline) 3.31 6.97
$\tau\sim\mathcal{U}(0,1)$ (default) 2.77 5.56
$\mathrm{logit}(\tau)\sim\mathcal{N}(-0.8,1)$ 2.79 6.04
$\mathrm{logit}(\tau)\sim\mathcal{N}(0,1)$ 2.77 5.93
$\mathrm{logit}(\tau)\sim\mathcal{N}(0.8,1)$ 2.58 5.44
Table 1: Impact of noise level sampling.
Noise schedule.
Inspired by Hoogeboom et al. (2023); Esser et al. (2024), we study how the sampling distribution of the noise level $\tau$ in Eq. 4 affects generation quality. Following Esser et al. (2024), we draw $\tau$ from a logit-normal distribution with mean $\mu$ (and fix $\gamma=3.0$). As shown in Table 1, all uniform and non-uniform schedules improve FID over the $\tau=0$ baseline, and concentrating more probability mass on higher noise levels ($\mu=0.8$) yields the relatively best results for both MAR-B and SiT-B. This indicates that $l$-DeTok is robust to the choice of noise schedule and can still benefit from further tuning of the noise distribution.
@@ 5.1.2 Properties of Masking
Masking ratio. We examine how varying the maximal masking ratio $M$ (Eq. 5) influences generation quality. As shown in Figure 3-(a), both SiT and MAR benefit from masking-based tokenizer training in generation quality. Masking ratios between 70% and 90% consistently yield stronger performance compared to low masking ratios (e.g., 30%), favoring high degrees of masking. This behavior mirrors observations from latent denoising, i.e., challenging denoising is more beneficial, indicating a common underlying principle under the deconstruction-reconstruction strategy.
Constant vs. randomized masking ratio. Figure 3-(b) compares randomized masking ratios against constant ones. For constant ratios, we further fine-tune the tokenizer decoder for 10 epochs using fully-visible latents to alleviate the distribution shift between training and inference since fully-visible inputs are absent in constant-ratio training. This adjustment improves rPSNR by 1–3 and rFID by 0.6 (details in Sec. B.1). Figure 3-(b) shows randomized masking consistently outperforms constant masking, as it encourages latent embeddings to be robust to varying corruption levels. This aligns naturally with downstream tasks that require denoising across diverse corruption levels.
@@ 5.1.3 Joint Denoising
Prior ablations indicate that both latent noising and masking can independently improve generation quality, with latent noising showing a stronger effect. Here, we investigate the effect of joint denoising. Based on prior results, we fix the noise standard deviation to $\gamma=3.0$ and the masking ratio to $M=0.7$. Figure 5 and Table 5 summarize the results. With joint denoising, our $l$-DeTok achieves FID scores of 5.50 (SiT-B) and 2.65 (MAR-B) with CFG (Table 5). In comparison, our baseline tokenizer—trained with identical settings but without any noise—obtains worse results: 6.97 (SiT-B) and 3.31 (MAR-B). We observe that joint denoising is more effective for MAR, but provides limited additional benefit for SiT when latent noising is already applied. This indicates that latent denoising is essential, while the masking-based denoising is optional.
Lastly, with joint denoising, we increase the encoder to base size, train for 200 epochs, and enable the GAN loss starting from epoch 100 (+Extended in Table 5). Under this setting, FID improves to 5.13 (SiT) and 2.43 (MAR). We adopt this improved tokenizer for all subsequent evaluations.
Figure 4: Impact of tokenizer training strategies on generative performance. We compare our baseline tokenizer with our $l$-DeTok variants: masking-only (M=0.7M{=}0.7), latent noising-only ($\gamma{=}3.0$), and their combination (joint noising). Both masking and latent noising independently improve generation quality, with latent noising showing a stronger effect. Joint noising further improves performance for MAR, particularly in inception scores (IS), but provides limited additional benefit for SiT when latent noising is already applied. FID@50k scores are detailed in Tab. 5.
@@
[…]
Baseline Ours
SiT-B 7.08 5.13
SiT-L 4.66 3.49
SiT-XL 4.47 3.14
MAR-B 3.65 2.43
MAR-L 2.44 2.08
Table 2: Scalability.
5.1.4 Scalability
We study whether the gains from $l$-DeTok persist under generative model scaling. We train SiT-B/L/XL and MAR-B/L for 100 epochs, comparing the baseline tokenizer and ours under identical settings. As shown in Table 2, $l$-DeTok consistently improves FID at all model sizes for both SiT and MAR. The advantage remains when we extend training to 800 epochs in the system-level comparison of Table 4.
@@ 5.2 Generalization Experiments
To comprehensively evaluate tokenizer generalizability, we compare performance across six representative generative models: three non-autoregressive (DiT (Peebles and Xie, 2023), SiT (Ma et al., 2024), LightningDiT (Yao and Wang, 2025)) and three autoregressive (MAR (Li et al., 2024a), RandomAR, RasterAR). RandomAR and RasterAR are adapted from RAR (Yu et al., 2024). We modify them to support decoding continuous tokens via diffloss MLPs (Li et al., 2024a). RandomAR generates tokens in a random order while RasterAR uses a raster-scan order. We use base-sized models and train them for 100 epochs for experiments here.
Comparisons with standard convolutional tokenizers. Our $l$-DeTok tokenizer consistently outperforms conventional tokenizers (Rombach et al., 2022; Esser et al., 2021; 2024), by large margins across both AR and non-AR models. Table 3 presents the results. Compared to the best existing tokenizer (MAR-VAE), our method significantly improves FID (with CFG) from 3.71 to 2.43 ($\scriptstyle\sim$34%) for MAR, from 11.78 to 5.22 ($\scriptstyle\sim$56%) for RandomAR, and from 7.99 to 4.46 ($\scriptstyle\sim$44%) for RasterAR. Improvements for non-autoregressive models are also consistent.
Comparisons with semantics-distilled tokenizers. Our $l$-DeTok generalizes significantly better than prior semantics-distilled tokenizers. Table 3 compares our method with recent approaches such as VA-VAE (Yao and Wang, 2025) and MAETok (Chen et al., 2025a), which distill semantics from pretrained encoders. Surprisingly, we empirically find these tokenizers, despite promising performance in non-AR models, do not generalize well to AR models. Previous studies implicitly assume that tokenizer improvements from one generative paradigm naturally transfer to others. However, our experiments challenge this assumption, revealing a previously unrecognized gap: tokenizer effectiveness in one paradigm does not necessarily transfer to others.
In sharp contrast, our method generalizes robustly across both non-AR and AR models. Critically, our $l$-DeTok achieves this without any semantics distillation from pretrained encoders. To further investigate whether our $l$-DeTok can benefit from semantics distillation, we incorporate an auxiliary semantics-distillation loss (details in Sec. A.4). Remarkably, this privileged version of our tokenizer achieves the best FID scores for non-AR models, surpassing all previous semantics-distilled tokenizers. Although AR models benefit relatively less from distillation and may even slightly degrade, their performance still remains substantially superior to previous methods.
$l$-DeTok with convolutional tokenizers. Our $l$-DeTok generalizes across Transformer- and convolution-based tokenizers, as the idea of latent denoising is architecture-agnostic. To demonstrate this, we train CNN-based tokenizers for 50 epochs with and without our denoising objective (details in Sec. A.5). With the default CNN baseline, MAR-B achieves 3.32 FID and SiT-B 7.11 FID at best CFGs. With our denoising-based CNN tokenizers, MAR-B improves to 2.82 FID and SiT-B to 5.62 FID, demonstrating that $l$-DeTok provides consistent gains also in convolutional networks.
@@
VA-VAE (Yao and Wang, 2025) 0.28 16.66 144.5 38.13 68.3 15.88 160.5 4.33 221.1 4.91 213.9 2.86 275.1
MAETok (Chen et al., 2025a) 0.48 6.99 201.8 24.83 97.6 15.92 127.2 4.77 243.2 5.24 224.7 3.92 273.3
Our $l$-DeTok + Distillation 0.85 2.52 254.1 5.57 180.3 11.99 158.8 3.40 232.6 3.91 221.7 2.18 243.0
@@ 5.3 Benchmarking with Previous Systems
Table 4: System-level comparison on ImageNet $\times256$ class-conditioned generation. Our approach enables MAR models (Li et al., 2024a) to achieve leading results without relying on semantics distillation. : With additional decoder fine-tuning (see Sec. A.3 for details).
@@
SiT-XL + REPA (Yu et al., 2025b) 675M 5.90 157.8 0.70 0.69 1.42 305.7 0.80 0.64
SiT-XL + MAETok (Chen et al., 2025a) 675M 2.31 216.5 - - 1.67 311.2 - -
LightningDiT + MAETok (Chen et al., 2025a) 675M 2.21 208.3 - - 1.73 308.4 - -
LightningDiT + VAVAE (Yao and Wang, 2025) 675M 2.17 205.6 0.77 0.65 1.35 295.3 0.79 0.65
DDT-XL (Wang et al., 2025) 675M 6.27 154.7 0.68 0.69 1.26 310.6 0.79 0.65
@@
DiT-XL/2 (Peebles and Xie, 2023) 675M 9.62 121.5 0.67 0.67 2.27 278.2 0.83 0.57
SiT-XL/2 (Ma et al., 2024) 675M 8.30 - - - 2.06 270.3 0.82 0.59
VAR-d30 (Tian et al., 2025) 2.0B - - - - 1.92 323.1 0.82 0.59
LlamaGen-3B (Sun et al., 2024) 3.1B - - - - 2.18 263.3 0.81 0.58
RandAR-XXL (Pang et al., 2025) 1.4B - - - - 2.15 322.0 0.79 0.62
CausalFusion (Deng et al., 2024) 676M 3.61 180.9 0.75 0.66 1.77 282.3 0.82 0.61
MAR-B + MAR-VAE (Li et al., 2024a) 208M 3.48 192.4 0.78 0.58 2.31 281.7 0.82 0.57
MAR-L + MAR-VAE (Li et al., 2024a) 479M 2.60 221.4 0.79 0.60 1.78 296.0 0.81 0.60
MAR-H + MAR-VAE (Li et al., 2024a) 943M 2.35 227.8 0.79 0.62 1.55 303.7 0.81 0.62
MAR-B (Li et al., 2024a) + our $l$-DeTok 208M 2.79 195.9 0.80 0.60 1.61 289.7 0.81 0.62
MAR-B (Li et al., 2024a) + our l-DeTok 208M 2.94 195.5 0.80 0.59 1.55 291.0 0.81 0.62
MAR-L (Li et al., 2024a) + our $l$-DeTok 479M 1.84 238.4 0.82 0.60 1.43 303.5 0.82 0.61
MAR-L (Li et al., 2024a) + our l-DeTok 479M 1.86 238.6 0.82 0.61 1.35 304.1 0.81 0.62
[…]
ADM (Dhariwal and Nichol, 2021) 554M 7.72 172.7
DiT-XL/2 (Peebles and Xie, 2023) 675M 3.04 240.8
SiT-XL/2 (Ma et al., 2024) 675M 2.62 252.2
SiT-XL/2 + REPA (Yu et al., 2025b) 675M 2.08 274.6
MAR-L + MAR-VAE (Li et al., 2024a) 479M 1.73 279.9
MAR-B + Ours (scratch, 400ep) 208M 1.83 279.6
MAR-L + Ours (fine-tune, 200ep) 479M 1.61 315.7
Table 5: System-level comparison on ImageNet $\times512$ class-conditioned generation. Our tokenizer is fine-tuned from ImageNet $\times256$ checkpoint for 25 epochs. Scratch vs. fine-tune: Generative models trained from scratch or initialized from ImageNet $\times256$ checkpoints (see Sec. A.3 for details).
5.4 Text-to-Image Generation
We further validate $l$-DeTok on text-to-image (T2I) generation. Following the protocol in Bao et al. (2022); Yu et al. (2025b), we train models from scratch on the MS-COCO train split (Lin et al., 2014) and evaluate on the validation split using FID-30k. We compare T2I variants of MAR-B and SiT-B under identical conditions, varying only the tokenizer (details in Sec. C). Table 6 summarizes the results, demonstrating that $l$-DeTok substantially improves not only sample quality and diversity (FID), but also text–image conditioning alignment (CLIP score). Qualitatively (see Figures. C.1, C.2,C.3, and C.4), other tokenizers frequently produce the “spot artifacts” reported in previous works (Fan et al., 2025; Team et al., 2025) under the text-to-image setting, whereas such artifacts are notably absent with $l$-DeTok.
Tokenizer T2I MAR-B T2I SiT-B
[…]
VA-VAE (Yao and Wang, 2025) 34.64 21.98 5.83 25.07
SD-VAE Rombach et al. (2022) 19.75 22.86 6.63 24.00
MAR-VAE (Li et al., 2024a) 12.49 22.07 5.74 23.50
Ours ($l$-DeTok) 4.97 24.82 4.31 24.61
Table 6: MS-COCO text-to-image generation. Comparison of tokenizers on T2I-variants MAR-B and SiT-B, reported with best CFG scales. Our $l$-DeTok achieves both lower FID (better diversity) and higher CLIP scores (better alignment), outperforming all others.
5.5 Beyond Standard Continuous Tokenizers
To assess the generality of our approach beyond standard 2D continuous tokenizers, we study both 1D continuous tokenizers and discrete vector-qunatized (VQ) tokenizers.
1D Tokenizers.
We first test $l$-DeTok with recent 1D tokenizers that map images to compact 1D token sequences (Yu et al., 2025a; Yan et al., 2025; Duggal et al., 2025; Bachmann et al., 2025). These tokenizers substantially reduce the sequence length compared to standard 2D latent grids. Sec. A.6 provides implementation details and an interesting pitfall coupled with the “grokking” phenomenon we observe in these 1D tokenizers. As shown in Table 8, $l$-DeTok improves generation performance over the corresponding baselines, with gains becoming more pronounced at higher token budgets. This suggests that the benefit of denoising is not tied to a particular latent topology, and remains effective even when the representation is compressed and less-structured.
Vector-quantized Tokenizers.
We next evaluate $l$-DeTok on VQ tokenizers, where the latents are discrete code indices instead of continuous features. In this setting, the idea of removing masking noise applies directly. Section A.7 presents implementation details. We consider both random-order (RandomAR) and standard raster-scan auto-regressive models (RasterAR). As summarized in Table 8, $l$-DeTok improves both variants: for RandomAR, FID is improved from 8.48 to 7.03; for RasterAR, FID improves from 5.29 to 4.91, with IS also increasing in both cases. These gains indicate the broader applicability of our denoising principle beyond continuous tokenizers.
Table 7: 1D continuous tokenizers on ImageNet $\times256$. Comparison of MAR and SiT with different numbers of 1D tokens, with and without our denoising training, evaluated by FID@50k.
MAR-B SiT-B
#Tokens Setting w/ CFG w/o CFG w/ CFG w/o CFG
32 Baseline 5.22 17.63 6.99 18.91
  +Ours 5.51 14.87 7.02 16.46
64 Baseline 4.58 17.73 7.05 22.82
  +Ours 4.12 14.02 6.40 18.59
128 Baseline 3.92 18.63 7.95 28.36
  +Ours 3.14 12.05 6.20 20.22
Table 8: Vector-quantized tokenizers on ImageNet $\times256$. Comparison of VQ tokenizers with and without our denoising training in RandomAR and RasterAR models, evaluated by FID@50k and IS@50k with CFGs. $l$-DeTok improves generative quality across generation orders.
Setting RandomAR-B RasterAR-B
[…]
Baseline VQ 8.48 144.93 5.29 204.15
  +Ours 7.03 162.68 4.91 208.61